OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 3x − 2y = 6.

OpenStudy (anonymous):

x-intercept: (0, −3), y-intercept: (0, 2) x-intercept: (−3, 0), y-intercept: (2, 0) x-intercept: (0, 2), y-intercept: (0, −3) x-intercept: (2, 0), y-intercept: (0, −3)

OpenStudy (amistre64):

\[ax+by=c\] \[\frac{ax}{c}+\frac{by}{c}=\frac{c}{c}\] \[\frac{x}{c/a}+\frac{y}{c/b}=1\] xint = c/a yint = c/b

OpenStudy (amistre64):

looks to be the last one

OpenStudy (anonymous):

you can also do this by setting x to zero and solving for y. and vice versa to find x-intercept.